OpenStudy (anonymous):

Please help its my last assignment and I'm desperate A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t)=-16t^2+34t+80 The average rate of change of f(t) from t=5 seconds to t=7 seconds is _____ feet per second.

OpenStudy (bibby):

\(\huge A(x)=\frac{f(x)-f(a)}{x-a}\) A being the name of the average rate of change function

OpenStudy (anonymous):

So whats the answer?

OpenStudy (bibby):

Plug in and find out what f(7) and f(5) are

OpenStudy (anonymous):

how

OpenStudy (bibby):

f(t) = \(\large -16t^2+34t+80\) f(7) = \(\large -16(7^2)+34(7)+80\)

OpenStudy (anonymous):

so confused I'm sorry

OpenStudy (bibby):

I'll write it out

OpenStudy (bibby):

if we wanted the average rate of change from t=3 to t = 5 we would use the following: \(\huge A(x)=\frac{f(5)-f(3)}{5-3}\) now what is f(5)? f(5) = \(\large -16(5^2)+34(5)+80)\)

OpenStudy (anonymous):

i only have a min to submit :(

OpenStudy (bibby):

I'm not going to give you the answer :(

OpenStudy (anonymous):

k thanks anyways

OpenStudy (bibby):

שלום

OpenStudy (bibby):

;~; שלום http://puu.sh/b9csX/2e982ab6ad.png

OpenStudy (anonymous):

-582?

OpenStudy (bibby):

−16(7^2)+34(7)+80 -16*49=-784+238+80=-466 just plug it into a calculator... f(7) = -466

OpenStudy (anonymous):

i did thats what i got

OpenStudy (bibby):

I have no idea what you're even doing right now. find f(7) and f(5)